# NetMamba: Efficient Network Traffic Classification via Pre-training Unidirectional Mamba

### v3

@@ VII-D2 Data-level Ablation
When all header bytes in a packet are omitted, classification performance declines significantly, with accuracy dropping by 14.51% to 48.54%. This highlights the critical role of key fields within packet headers—such as port number, protocol, and packet length—which have been proven effective in traffic classification [4, 38, 39].
Regarding packet payloads, the ablation results show accuracy drops ranging from 0.26% to 8.29% across five datasets. This underscores the contribution of potential plaintext and specific encrypted payloads to improved traffic understanding.
Without IP masking, the model may learn biased shortcuts based on IP addresses present in the training set, resulting in a maximum decrease in accuracy of 6.27% in the test set.
Likewise, the vertical bias information introduced by the 2-dimensional patch splitting results in a maximum accuracy decline of 0.97%, highlighting the importance of 1-dimensional stride cutting.
@@ VII-E Few-Shot Evaluation
Among the pre-trained methods, ET-BERT shows less reliability across the four datasets. YaTC performs similarly to NetMamba on three datasets for most labeled data sizes but falls significantly behind NetMamba on the CrossPlatform (Android) dataset. Therefore, our model exhibits excellent robustness and is highly effective at solving classification problems with limited encrypted traffic data.

### v4

@@ VII-D2 Data-level Ablation
When all header bytes in a packet are omitted, classification performance declines significantly, with accuracy dropping by 15.51% to 48.75%. This highlights the critical role of key fields within packet headers—such as port number, protocol, and packet length—which have been proven effective in traffic classification [4, 38, 39].
Regarding packet payloads, the ablation results show accuracy drops ranging from 0.16% to 0.84% across four datasets. This underscores the contribution of potential plaintext and specific encrypted payloads to improved traffic understanding.
Likewise, the vertical bias information introduced by the 2-dimensional patch splitting results in a maximum accuracy decline of 1.88%, highlighting the importance of 1-dimensional stride cutting.
@@ VII-E Few-Shot Evaluation
Among the pre-trained methods, ET-BERT demonstrates lower reliability on two datasets, while YaTC performs comparably to NetMamba on all datasets. Consequently, our model exhibits exceptional robustness, on par with Transformer-based models, and proves highly effective in addressing classification tasks with limited encrypted traffic data.